OpenStudy (shamil98):

I need some help with eccentricity. The question will be posted in the comments.. I know e = c/a How do I find c?

OpenStudy (shamil98):

\[\large \frac{ (x-5)^2 }{ 8 } + \frac{ (y+1)^2 }{ 12 } = 1\]

OpenStudy (shamil98):

b^2 + c^2 = a^2 a^2 = 12 a = 2√3 b^2 = 8 b = 2√2 8 + c^2 = 12 c^2 = 4 c = 2 so.. e = 2/2√3 e = 1/√3 right?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

a = major axis b = minor axis so looks in ur calucaltion, u have reversed them

ganeshie8 (ganeshie8):

oh no, you're right ! i overlooked it ...

OpenStudy (shamil98):

the major axis is equal to 2a. minor axis is equal to 2b. and a^2 is always the bigger one of the two denominators.. atleast that's what i read ..

ganeshie8 (ganeshie8):

yes ! looks good

OpenStudy (shamil98):

thanks.

ganeshie8 (ganeshie8):

np :)